OpenStudy (anonymous):

Sum of the series 1 + 2.2 + 3.2^2 + 4.2^3.....+ 100.2^99 is?

OpenStudy (anonymous):

\[a_n=n\times2^{n-1}\]

OpenStudy (anonymous):

gimmick is this is the derivative

OpenStudy (anonymous):

The answer should be 99.2^100 + 1

OpenStudy (anonymous):

i take it these are multiplication signs, not decimal points right?

OpenStudy (anonymous):

Aha..u r point there @satellite73

OpenStudy (anonymous):

\[1+r+r^2+r^2+...+r^n=\frac{r^{n+1}-1}{r-1}\]

OpenStudy (anonymous):

take the derivative of both sides and evaluate at \(r=2\)

OpenStudy (anonymous):

and it is AGP

OpenStudy (anonymous):

@satellite73 @amistre64 @mukushla r u there

OpenStudy (anonymous):

lol...everyone is here

OpenStudy (anonymous):

Come...On......PLzz help

OpenStudy (anonymous):

man satellite mentioned all the things u need to solve

OpenStudy (anonymous):

i Did nt understand..

OpenStudy (anonymous):

I think it is AGP..!

OpenStudy (anonymous):

So.. it is Very Difficult to solve by that.....IF u have a easy method..plzz Tell me

OpenStudy (anonymous):

oh its very easy man...can u take derivative?

OpenStudy (anonymous):

Yup.. i knw to differentiate...

OpenStudy (anonymous):

but did nt understand..satterlite

OpenStudy (anonymous):

\[1+r+r^2+r^2+...+r^{100}=\frac{r^{101}-1}{r-1}\]take derivative of both sides u will get: \[1+2r+3r^2+4r^2+...+100r^{99}=\frac{101r^{100}(r-1)-(r^{101}-1)}{(r-1)^2}\]now just let \(r=2\)

OpenStudy (anonymous):

101.2^100 - 2^101 + 1

OpenStudy (anonymous):

yeah and we are done

OpenStudy (anonymous):

but should it be 101.2^100 - 2^100 + 1

OpenStudy (anonymous):

yup..i got it,,,))

OpenStudy (anonymous):

man be carefull... \[101\times2^{100} - 2^{101} + 1=101\times2^{100} -2\times 2^{100} + 1=99\times2^{100}+ 1\]

OpenStudy (anonymous):

thxx..... Can u tell me Why we r differen it.??

OpenStudy (anonymous):

i only knw some basics on Calculus

OpenStudy (anonymous):

because we want to change it to the form that we need

OpenStudy (raden):

any else ...i think more eassy Let P = 1 + 2.2 + 3.2^2 + 4.2^3.....+ 100.2^99 times 2 to both side, thus we get : 2P = 2 + 2.2^2 + 3.2^3 + 4.2^4.....+ 100.2^100 2P - P = - 1 - 2 - 2^2 - 2^3 - 2^4 - ... - 2^99 + 100*2^100 P = - (1+2+2^2+2^3+2^4+...+2^99) + 100*2^100 P = - (1(2^100 - 1)/(2-1)) +100*2^100 P = - 2^100 + 1 + 100*2^100 P = 99*2^100 + 1

OpenStudy (anonymous):

thats really easier and better @RadEn thank u

OpenStudy (raden):

yw) but i need long time to solve it.... :D